# Embodied AI Agents: Modeling the World

### v1

@@ Abstract
This paper describes our research on AI agents embodied in visual, virtual or physical forms, enabling them to interact with both users and their environments. These agents, which include virtual avatars, wearable devices, and robots, are designed to perceive, learn and act within their surroundings, which makes them more similar to how humans learn and interact with the environments as compared to disembodied agents. We propose that the development of world models is central to reasoning and planning of embodied AI agents, allowing these agents to understand and predict their environment, to understand user intentions and social contexts, thereby enhancing their ability to perform complex tasks autonomously. World modeling encompasses the integration of multimodal perception, planning through reasoning for action and control, and memory to create a comprehensive understanding of the physical world. Beyond the physical world, we also propose to learn the mental world model of users to enable better human-agent collaboration.
@@ 5.3.1 Goal Inference Benchmark
Figure 10: A multi-modal sample from the benchmark. Here the video and digital contexts are relevant to the prediction problem, and audio/longitudinal are noise.
@@ 5.3.2 WorldPrediction Benchmark
We refer to the underlying details of WorldPrediction in section 3.6.4. This benchmark also serves as a practical touchstone for the specific challenges faced by wearable agents. Egocentric devices must transform a rapidly shifting first-person video stream into compact, predictive representations that guide on-device assistance (e.g., suggesting the next tool while cooking or flagging a safety hazard before it occurs). Because its two tasks require choosing high-level actions and full action sequences under partial observability, WorldPrediction mirrors the temporal abstraction and uncertainty inherent to on-body perception. Moreover, its “action-equivalent” design forces models to rely on inferred causal structure rather than pixel-level continuity—exactly the type of abstraction a resource-constrained wearable must compute to remain responsive. Thus, evaluating an agent’s policy generator on WorldPrediction yields a direct measure of whether it can anticipate user intent several steps ahead, a prerequisite for proactive, context-aware assistance in open-world settings.
@@ 6.1.1 Physical Capabilities
Manipulation involves any purposeful interaction with objects such as generalized grasping and placing of objects, as well as advanced and meaningful tool use. While basic capabilities of grasping (and to some extend) placing have made significant progress, we’re yet to see an agent that can perform truly general pick&place. An important sub-category of manipulation is dexterous manipulation, which involves finegrained control of objects (e.g., “put the key in the lock”, "re-orient object in hand"), often requiring or benefiting from using multi-fingered hands. This area is even more challenging than the more coarse picking and placing manipulation skills, and is still an open research problem.
While most research focuses on one of the physical capabilities, there is an increased focus on learning policies for "full body control" which attempts to learn policies that can control both the location of the robot in it’s environment as well as the manipulation of the environment.
6.1.2 "Brain" Capabilities
@@
A robotic agent must be capable of following language instructions at multiple levels of abstractions and must be able to break down such task requests into smaller units of activities. For instance, when asked to perform the task of "clean up the kitchen", the agent needs to semantically understand the request, figure out what sub tasks to perform ("find dirty dishes", "bring dirty dishes to dishwasher", "empty dishwasher first if needed", …) and then plan both at that abstract task level as well as at the more concrete sub-task level in dynamic environments.
@@ 6.2 Architecture and Models
Both of these approaches inherently assume that generalization emerges as a function of data scale (whether teleoperated or through interaction in the simulator) and that a robot agent can "learn once everything the robot needs to know". However it is entirely unclear whether it is feasible that we can prepare the robot for everything it might encounter in the real world, or whether we can even learn a policy that can represent all possible skills in every possible scenario. Instead of trying to do so, we should also enable the robot to understand how its actions affect the world (through learning a world model), such that it can find new solutions at test (inference time). World models are a promising route to achieving general robot behaviors with limited in-domain data, and naturally enable training from various (heterogeenous) data sources (successful and unsuccessful task executions, video data, play data). Ultimately we believe that a system combining policies, world models and "reward" models will be our best bet to general purpose robotic agents (see Figure 11 (left)).
@@ 6.2.1 A Note on Classical Robotics utilizing Analytical Models
To enable extract extractions from a world model we need to perform reasoning/planning with a world model, and we imagine to do this similar to how model predictive control is performed. Model predictive control with (analytical) world models has a long and rich history in robotics, especially for "lower" level control of robot motions (see Figure 11 for an example). Although here we discuss learned (action-conditioned) world models (and discuss world models at different levels of action abstractions), it is not necessary to reinvent the wheel on the axis of planning with world models. Instead we can leverage and/or build on advances in robotic control with analytical models.
@@ 6.3.2 Hardware Benchmarks
Due to the challenges of setting up and maintaining a robot evaluation environment, large-scale public robot benchmarks that evaluate generalist robots across diverse environments do not exist. However some explorations in this direction exist. Challenges like the HomeRobot Challenge (Yenamandra et al., 2023) provide a simulation platform for researchers to prototype and test their algorithms, and then invite best performers to also test their algorithm on hardware. There are also first attempts at providing an "evaluation on hardware service" that allows researchers to submit policies that then get evaluated either fully automatically through learned success classifiers (Zhou et al., 2025), or through deploying human evaluators that compare and rank policies (Atreya et al., 2025). These explorations are in its infancy, and many questions remain unanswered, such as how to robustly extract metrics in the non-stationary environment (robot and environments change over time) these evaluations are performed on.

### v2

@@ Abstract
This paper describes our research on AI agents embodied in visual, virtual or physical forms, enabling them to interact more naturally with both users and their environments. These agents, which include virtual avatars, wearable devices, and robots, are designed to perceive, learn and act within their surroundings, which makes them more similar to how humans learn and interact with the environments as compared to disembodied agents. We propose that the development of world models is central to reasoning and planning of embodied AI agents, allowing these agents to understand and predict their environment, to understand user intentions and social contexts, thereby enhancing their ability to perform complex tasks autonomously. World modeling encompasses the integration of multimodal perception, planning through reasoning for action and control, and memory to create a comprehensive understanding of the physical world. Beyond the physical world, we also propose to learn the mental world model of users to enable better human-agent collaboration.
@@ 3.6.4 World Prediction Benchmark
Figure 9: WorldPrediction benchmark evaluates the high-level world modeling and long-horizon procedural planning capabilities of models from observations only, where the actions are depicted as videos and the states as images.
@@ 4 Type I: Virtual Embodied Agents
More recently, Seamless Interaction (Agrawal et al., 2025) introduces a family of audiovisual dyadic motion models compatible with both 2D and 3D renderings. The models, conditioned on speech from two parties as the input, can jointly generate facial expressions and body gestures. The motion models produce gestures and expressions of one specific speaker while taking into consideration the audio from both people. This allows the models to visualize speaking gestures, listening gestures, and turn-taking cues. The AV Dyadic Motion models go one step further to show visual synchrony by also taking into consideration the visual input of the other party.
4.1 Capability Taxonomy
The virtual agent boasts a range of cutting-edge capabilities that enable it to deliver highly realistic and engaging interactions. One of its key capabilities to generate full-body movements, including facial expressions, arm and hand gestures. This allows the agent to convey emotions and intentions in a more nuanced and human-like way. Another significant feature of a virtual agent is its dyadic audiovisual conditioning capability. This means that the model can take into account speech from two parties, whether it is human-generated or produced by a speech large language model (aka, speech LLM). Additionally, the agent can optionally condition on visual input from users, enabling interesting visual synchrony phenomena such as smile mirroring. This feature enables the agent to respond more naturally and empathetically to user input. The virtual agent also offers face controllability, allowing models to be fine-tuned for greater expressiveness. This has significant potential applications in building more attentive and empathetic virtual listeners, which can be particularly valuable in fields like customer service and education. In terms of gesture generation, virtual agents could produce illustrative gestures that match the semantic content of the conversation. For example, when saying the word "fly," the agent will extend its arms to mimic the action of flying. This adds an extra layer of realism and engagement to the interaction. Besides motion generation, the research of virtual agents has made significant strides in integrating motion models with speech LLM which play a role of spoken dialogue generator. The output of a speech LLM can visually guide the motion of the agent, creating a more seamless and natural interaction. This integration of speech and motion models paves the way for even more sophisticated and intuitive interactions in the future. Finally, the output representations of motion models should be visualized by the virtual agent. For example, virtual agents could use 2D and 3D renderings to demonstrate visualizations, which is critical for a fully immersive VR experience.
4.2 Architecture and Models
(a) Cascaded integration.
(b) Codebook integration.
Figure 10: Architecture of our Dyadic Motion Model
The virtual agent comprises of two key components: a speech LLM which handles spoken dialogue generation, and motion model which generates visual behavior. Figure 10(a) shows a straightforward approach to the integration of motion model with speech model via the generated speech. A more advanced approach is the integration via codebook as in Figure 10(b). Adapters are built on top of a speech LLM, taking its hidden states as input and predicting emotion and gesture codes. These codes serve as visual guidance for motion models besides the dyadic speech. The codebook integration is motivated by speech LLM’s ability to grasp conversational context, enabling the extra controllability of generation of gestures and expressions that suit each moment of the conversation.
Figure 11: Dyadic Motion Model is conditioned on speech tokens obtained from dyadic audio and optionally user’s visual features, and generates face Imitator features and body SMPL-H features.
We then share details of dyadic motion models as part of the virtual agent. Dyadic Motion Model in Figure 11 is based on a Diffusion Transformer architecture trained with flow-matchin objective. The model is conditioned on dyadic audio input and optionally user’s visual features. There are controllable variants of motions models that can adapt emotion responses and expressivity levels, as well as generating more semantically-relevant gestures. As for output, the motion model generates Imitator features as face representations, and SMPL-H as body movement features. These face and body features are compatible with both 2D and 3D renderings.
4.3 Benchmarks
The goal of virtual agents is to create realistic, engaging, and effective human-computer interactions. The key research topics are to design and evaluate virtual agents that can simulate human-like behavior, conveying emotions, intentions, and empathy through not only speech, but also facial expressions and body language.
Dataset.
Seamless Interaction Dataset is vast and diverse collection of in-person, dyadic interactions designed to drive innovation in social AI research. With over 4,000 hours of interactions from more than 4,000 participants, this dataset provides an unparalleled wealth of data for researchers to explore. The dataset features nearly 1,300 conversational and activity-based prompts, covering a wide range of topics, interpersonal stances, and participant relationships. Anchored in contemporary psychological theory, the dataset includes both Naturalistic and Improvised content, offering a rich and nuanced understanding of human interaction. To further contextualize the interactions, the dataset is accompanied by detailed annotations and metadata, providing researchers with a comprehensive foundation for developing and evaluating social AI models. By leveraging this extensive dataset, researchers can advance our understanding of human behavior and develop more sophisticated and effective social AI systems.
Subjective metrics.
Evaluating behavior in embodied agents remains an unsolved problem, Agrawal et al. (2025) presents an approach to human subjective studies that incorporate face- and body-dyadic protocols. The face protocol focuses on facial expressions and head movements in photorealistic renderings. As for body protocol, it pays attention to visual behavior involving overall pose, hands, arms, shoulders, and head, but without facial rendering. The human studies use a pairwise approach where annotators view two dyads side-by-side, and participants provided a preference rating for each pair of visualizations. The protocols consist of $10$ core evaluative dimensions, which are grouped into three sections: overall preferences, listener-behavior preferences, and speaking-behavior preferences. The evaluation dimensions include key aspects such as lifelikeness, clarity of intent, turn-taking, listening, and speaking. By examining these critical dimensions, we can gain a deeper understanding of how well virtual agents simulate human-like behavior and interact with humans in a natural and effective way.
Objective metrics.
The research community have adopted automatic metrics to quantitatively evaluate the visual quality of face and body generations. In terms of face metrics, Lip-Sync scores including Sync-C and Sync-D evaluate the lip synchronization with speech, and Fréchet Inception Distance (FID) to assess face image quality based on features encoded by the Inception network. For body generation, Fréchet Gesture Distance (FGD) is widely used to quantify the discrepancy in distribution between generated outputs and real data. Furthermore, Diversity is also used to measure the range of variations present in the generated gestures.
@@ 5.3.1 Goal Inference Benchmark
Figure 13: A multi-modal sample from the goal inference benchmark. Here the video and digital contexts are relevant to the prediction problem, and audio/longitudinal are noise.
@@ 5.3.2 WorldPrediction Benchmark
Figure 14: Observable vs. Occluded high-level states and actions in WorldPrediction. We show the environmental and body state changes between initial and final states taken from EgoExo4D (Grauman et al., 2024)
We refer to the underlying details of WorldPrediction in section 3.6.4. This benchmark also serves as a practical touchstone for the specific challenges faced by wearable agents. Egocentric devices must transform a rapidly shifting first-person video stream into predictive representations that guide on-device assistance (e.g., suggesting the next tool while cooking or flagging a safety hazard before it occurs). Because its two tasks require choosing high-level actions and full action sequences under partial observability, WorldPrediction mirrors the temporal abstraction and uncertainty inherent to on-body perception. Moreover, its “action-equivalent” design forces models to rely on inferred causal structure rather than pixel-level continuit, exactly the type of abstraction a resource-constrained wearable must compute to remain responsive. WorldPrediction also includes egocentric action segment with exocentric state observations from EgoExo4D as shown in figure 14. Thus, evaluating an agent’s policy generator on WorldPrediction yields a direct measure of whether it can anticipate user intent several steps ahead, a prerequisite for proactive, context-aware assistance in open-world settings.
@@ 6.1.1 Physical Capabilities
Manipulation involves any purposeful interaction with objects such as generalized grasping and placing of objects, as well as advanced and meaningful tool use. While basic capabilities of grasping (and to some extend) placing have made significant progress, we’re yet to see an agent that can perform truly general pick&place. An important sub-category of manipulation is dexterous manipulation, which involves finegrained control of objects (e.g., “put the key in the lock”, “re-orient object in hand”), often requiring or benefiting from using multi-fingered hands. This area is even more challenging than the more coarse picking and placing manipulation skills, and is still an open research problem.
While most research focuses on one of the physical capabilities, there is an increased focus on learning policies for “full body control” which attempts to learn policies that can control both the location of the robot in it’s environment as well as the manipulation of the environment.
6.1.2 “Brain” Capabilities
@@
A robotic agent must be capable of following language instructions at multiple levels of abstractions and must be able to break down such task requests into smaller units of activities. For instance, when asked to perform the task of “clean up the kitchen”, the agent needs to semantically understand the request, figure out what sub tasks to perform (“find dirty dishes”, “bring dirty dishes to dishwasher”, “empty dishwasher first if needed”, etc.) and then plan both at that abstract task level as well as at the more concrete sub-task level in dynamic environments.
@@ 6.2 Architecture and Models
Both of these approaches inherently assume that generalization emerges as a function of data scale (whether teleoperated or through interaction in the simulator) and that a robot agent can “learn once everything the robot needs to know”. However it is entirely unclear whether it is feasible that we can prepare the robot for everything it might encounter in the real world, or whether we can even learn a policy that can represent all possible skills in every possible scenario. Instead of trying to do so, we should also enable the robot to understand how its actions affect the world (through learning a world model), such that it can find new solutions at test (inference time). World models are a promising route to achieving general robot behaviors with limited in-domain data, and naturally enable training from various (heterogeenous) data sources (successful and unsuccessful task executions, video data, play data). Ultimately we believe that a system combining policies, world models and “reward” models will be our best bet to general purpose robotic agents (see Figure 15 (left)).
@@ 6.2.1 A Note on Classical Robotics utilizing Analytical Models
To enable extract extractions from a world model we need to perform reasoning/planning with a world model, and we imagine to do this similar to how model predictive control is performed. Model predictive control with (analytical) world models has a long and rich history in robotics, especially for “lower” level control of robot motions (see Figure 15 for an example). Although here we discuss learned (action-conditioned) world models (and discuss world models at different levels of action abstractions), it is not necessary to reinvent the wheel on the axis of planning with world models. Instead we can leverage and/or build on advances in robotic control with analytical models.
@@ 6.3.2 Hardware Benchmarks
Due to the challenges of setting up and maintaining a robot evaluation environment, large-scale public robot benchmarks that evaluate generalist robots across diverse environments do not exist. However some explorations in this direction exist. Challenges like the HomeRobot Challenge (Yenamandra et al., 2023) provide a simulation platform for researchers to prototype and test their algorithms, and then invite best performers to also test their algorithm on hardware. There are also first attempts at providing an “evaluation on hardware service” that allows researchers to submit policies that then get evaluated either fully automatically through learned success classifiers (Zhou et al., 2025), or through deploying human evaluators that compare and rank policies (Atreya et al., 2025). These explorations are in its infancy, and many questions remain unanswered, such as how to robustly extract metrics in the non-stationary environment (robot and environments change over time) these evaluations are performed on.